# Boosting Generic Semi-Supervised Medical Image Segmentation via Diverse Teaching and Label Propagation

### v1

@@ Abstract
Both limited annotation and domain shift are significant challenges frequently encountered in medical image segmentation, leading to derivative scenarios like semi-supervised medical (SSMIS), semi-supervised medical domain generalization (Semi-MDG) and unsupervised medical domain adaptation (UMDA). Conventional methods are generally tailored to specific tasks in isolation, the error accumulation hinders the effective utilization of unlabeled data and limits further improvements, resulting in suboptimal performance when these issues occur. In this paper, we aim to develop a generic framework that masters all three tasks. We found that the key to solving the problem lies in how to generate reliable pseudo labels for the unlabeled data in the presence of domain shift with labeled data and increasing the diversity of the model. To tackle this issue, we employ a Diverse Teaching and Label Propagation Network (DTLP-Net) to boosting the Generic Semi-Supervised Medical Image Segmentation. Our DTLP-Net involves a single student model and two diverse teacher models, which can generate reliable pseudo-labels for the student model. The first teacher model decouple the training process with labeled and unlabeled data, The second teacher is momentum-updated periodically, thus generating reliable yet divers pseudo-labels. To fully utilize the information within the data, we adopt inter-sample and intra-sample data augmentation to learn the global and local knowledge. In addition, to further capture the voxel-level correlations, we propose label propagation to enhance the model robust. We evaluate our proposed framework on five benchmark datasets for SSMIS, UMDA, and Semi-MDG tasks. The results showcase notable improvements compared to state-of-the-art methods across all five settings, indicating the potential of our framework to tackle more challenging SSL scenarios.
Index Terms:
Semi-supervised medical image segmentation, Unsupervised Domain Adaptation, Domain Generalization, Semi-supervised Learning
@@ I Introduction
Semi-supervised medical image segmentation (SSMIS) methods [1, 2, 3] have recently gained growing interest due to its ability to overcome the significant challenges posed by the time-consuming and labor-intensive manual labeling process within the medical field. However, most existing SSMIS methods assume that the labeled and unlearned data sets follow the same distribution. This assumption restricts the application of these methods in situations where domain shift occurs. As a result, derivative scenarios like unsupervised medical domain adaptation (UMDA) [4, 5, 6] and semi-supervised medical domain generalization (Semi-MDG) [7, 8] have emerged. In the UMDA scenario, the training data comes from two different domains, and labels in the target domain are not accessible. In the Semi-MDG scenario, the training data is collected from multiple domains, but only a small portion of these domains have labeled data. Moreover, the testing data comes from an unseen domain, which makes the evaluation setting more challenging and closer to real-world situations.
Previous methods usually are task-specific designed to adopt one of these problems independently. Existing methods adopted for SSL may fail to solve the UDA problems, and vice versa [9, 10]. Since all of these frameworks involve learning from both labeled and unlabeled data, it is logical to develop a unified generic semi-supervised medical image segmentation framework that effectively manages both scenarios. Recently, AIDE [9] introduces an annotation-efficient deep learning framework to address three challenges (semi-supervised learning (SSL), unsupervised domain adaptation (UDA), and noisy label learning (NLL), leaving the Semi-MDG unaddressed. Furthermore, A&D [10] introduced an aggregating and decoupling approach to solve these three challenging tasks simultaneously. While their A&D framework has achieved notable improvements in all of these three tasks, its advancements in tasks involving domain shift, e.g., UMDA and Semi-MDG, remain limited compared to prior art methods that are specifically designed for these tasks. Furthermore, when applied to UMDA and Semi-MDG tasks, other state-of-the-art (SOTA) SSMIS methods [11, 12] exhibited even inferior performance compared to A&D; see results in Tab. In addition, the potential of the data structure and voxel relationships has yet not been fully explored, leading to suboptimal performance, especially for the UMDA and Semi-MDG scenarios. In addition, the noisy pseudo labels generated by one teacher model are not reliable and lack diversity.
The main challenge in semi-supervised medical image segmentation lies in efficiently leveraging unlabeled images, while UMDA and Semi-MDG scenarios further require addressing the domain shift issues. All of these three tasks benefit from the reliable pseudo-label and model diversity [2]. However, pseudo-labels inevitably include noise; simply incorporating these predictions into model training is likely to trigger the confirmation bias problem [13], i.e., overfitting to the incorrect pseudo-labels [14], especially under the large diversity conditions. On the other hand, one potential solution to mitigate the adversarial influence of domain shift is to learn the domain-invariant features. A&D [10] adopts the Diffusion-based V-Net to capture the domain-invariant features, which is not enough to learn the common knowledge of different domains. To address this challenge, we propose a Diverse Teaching and Label Propagation Network (DTLP-Net) boost the Generic Semi-Supervised Medical Image Segmentation by facilitating the delivery of complementary knowledge across these two flows; see Fig. 1. The core idea is to design strategies which can boost all of these three tasks. First, we employ A&D [10] as our baseline, where a Diffusion encoder that extracts domain-invariant features by building a shared knowledge base from multiple domains, and three decoders,one Diffusion VNet (D-VNet) decoder and two additional vanilla VNet decoders, that decouple the training process with labeled and unlabeled data, thus avoiding over-fitting to labeled data, specific domains and classes. Second, for the unlabeled data flow, we design the diverse teacher models to generate high quality pseudo labels. The first teacher prediction comes from the other tow decoders predictions, and we further adopt the mean-teacher model to generate the predictions, and then we employ an entropy-based ensembling strategy to encourage the model to learn from both the consistent and conflicting predictions between the teachers. This cross-supervision encourages the different decoders to learn complementary features and correct each other’s mistakes, enhancing the overall performance on unlabeled data. Third , to capture the global and local structure of the data, we utilize cross-set Cut-Mix strategy and masked image modeling to achieve consistency learning, which can also benefit for learning domain-invariant features. Finally, to fully capture the voxel correlations within the image, we propose label propagation to enhance the model’s robustness. With these two knowledge delivery mechanisms, a naive pseudo-labeling framework could handle the core challenge of generic SSMIS and surpass the SOTA methods designed for specific scenarios. The overall contributions can be summarized as follows:
@@
  • We propose an efficient Diverse Teaching and Label Propagation Network (DTLP-Net) to boost the Generic Semi-Supervised Medical Image Segmentation performance, including SSMIS, UMDA, and Semi-MDG tasks.
  • We design dual teacher model to enlarge the diversity of the two teacher models and an entropy-based ensembling strategy to encourage the model to learn from both the consistent and conflicting predictions between the teachers.
  • We propose cross-set Cut-Mix strategy and masked image modeling to to capture the global and local structure of the data, which can also benefit for learning domain-invariant features. Finally, to fully capture the voxel correlations within the image, we propose label propagation to enhance the model’s robustness.
@@
  • We evaluate our approach on five standard benchmark datasets, the experimental results show that our approach attains substantial enhancements across all three scenarios of SSMIS, UMDA, and Semi-MDG tasks.
@@ II-A Semi-supervised medical image segmentation
In an effort to reduce the costs related to volume-level annotation while still ensuring high accuracy, diverse semi-supervised methods for medical image segmentation tasks have been developed. Drawing inspiration from semi-supervised image classification [15], techniques such as self-training and consistency regularization. In addition, the mean-teacher approach [16] and its extensions [17, 18, 19] have gained significant attention. For instance, BCP [18] advocates a straightforward Mean Teacher architecture to copy and paste annotated and unlabeled data bidirectionally, highlighting a key challenge in semi-supervised learning where the distributions of labeled and unlabeled data may diverge. OTCMC [19] puts forward the utilization of Optimal Transport for the purpose of attaining efficient label propagation. Additionally, it employs Central Moment Consistency Regularization to direct the network’s focus towards the geometric structure of images.
Moreover, some works resort to the co-training paradigm, where multiple models are collaboratively trained by cooperatively training multiple models. CTCL [3] proposes a Collaborative transformer-CNN learning for semi-supervised medical image segmentation. MC-Net+ [20] addresses inconsistency through mutual consistency. W2sPC [21] further incorporates weak-to-strong perturbation consistency and edge-aware contrastive representation, thus promoting the consistency of reliable regions among diverse predictions and facilitating the learning of class-discriminative representations. GapMatch [22] effectively links instance and model perturbations, thereby expanding the perturbation space. Moreover, it utilizes dual perturbation to enforce consistency regularization on the model. Different these works that only consider one or two perturbations to enhance the model diversity, CMMT-Net [2] proposes a unified framework to generate reliable pseudo-labels and achieve consistency learning.
In contrast to the aforementioned techniques, we adopt co-training by multi-decoder to enlarge the diversity of the models, while design multi-teachers to enhance the pseudo-labels quality. This simple yet effective approach improves the model’s generalization ability and segmentation accuracy.
@@ II-B Unsupervised Medical Domain Adaptation
Unsupervised Domain Adaptation (UDA) [4, 23, 6] aims to solve the domain shifts by jointly training the model with labeled source domain data and unlabeled target domain data. In the field of Unsupervised medical Domain Adaptation (UMDA), one modality usually lacks any segmentation annotation, which has recently gained increasing attention since it offers an efficient way to compensate for limited medical image data. Existing UMDA approaches resort to various directions, e.g. semi-supervised learning [24], generative adversarial-based methods [25, 26, 27, 28], and contrastive learning [29], aiming to mitigate the adversarial influence caused by the severe domain shifts via image-level [4], feature-level [28] or both [23, 26, 5].
Though with promising adaptation results, these methods overlook the labeled source domain information and do not fully explore the unlabeled target domain information, hindering the generalization. In addition, the output-level alignment of different domains remains to explore.
@@ II-C Semi-supervised Medical Domain Generalization
Semi-supervised domain generalization (SemiDG) is a more challenging setting, where the model does not use any information from the target domain, and the data in the source domains are only partially labeled. Compared with SSMIS and UMDA tasks, SemiDG necessitates the model to possess robust feature extraction capabilities and a high degree of generalizability. Existing Semi-supervised domain generalization methods [30, 31] use various carefully designed strategies to solve the domain shifts, e.g., meta-learning [30], and Fourier transformation [7], which are not general and have unsatisfactory performance on the tasks such as UDA and SSL.
Considering all of these three tasks involve learning from both labeled and unlabeled data, A&D [10] first try to unified with a generic framework handling all of them. However, despite the advancement they achieved, the improvement on SSMIS, UMDA and Semi-MDG tasks is marginal compared with task-tailored SOTAs.
@@ III Methodology
In the 3D Generic Semi-Supervised segmentation setting, the training data comprises $N_{L}$ labeled samples $D_{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{N_{L}}$ and $N_{U}$ unlabeled samples $D_{u}=\left\{x_{i}\right\}_{i=1}^{N_{U}}$, respectively, where $x_{i}\subset\mathbb{R}^{L\times W\times H}$ depicts the input volume with dimensions $L\times W\times H$, and $y_{i}\subset\left\{0,1\right\}^{Y\times L\times H\times W}$ represents the ground truth with $Y$ classes. The aim is to utilize the datasets $D_{l}$ and $D_{u}$ to train a model to generate meaningful semantic predictions. Our DTLP-Net consists of a shared encoder $\mathcal{E}(x^{l},x^{u};\xi)$ to learn domain-invariant features, and three decoders, e.g., V-Net Decoder $\mathcal{D}(x^{l};\psi)$, Diffusion Decoder $\mathcal{D}(x^{l},x^{u};\xi)$ and V-Net Decoder $\mathcal{D}(x^{u};\theta)$, where $\psi$, $\xi$, and $\theta$ are model parameters. In addition, for $\mathcal{D}(x^{u};\theta)$ Decoder branch, we further obtain the mean-teacher Encoder $\mathcal{E}(x^{l},x^{u};\xi^{T})$ and Decoder $\mathcal{D}(x^{u};\theta^{T})$.
@@ III-A Preliminary baseline
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\epsilon\in\mathcal{N}(0,1)$ (1)
@@
$\mathcal{L}_{deno}=\frac{1}{N_{L}}\sum_{i=0}^{N_{L}}\mathcal{L}_{DiceCE}(p^{l;\xi}_{i},y_{i}),$ (2)
where $\mathcal{L}_{DiceCE}(x,y)=\frac{1}{2}[\mathcal{L}_{CE}(x,y)+\mathcal{L}_{Dice}(x,y)]$ is the combined dice and cross entropy loss.
Supervised Difficulty-aware Training with $D(x^{l};\psi)$
For the difficulty-aware training flow, i.e., $D(x^{l};\psi)$ as decoder, the encoder only takes $x_{l}$ input to obtain the multi-scale features $h^{l;\psi}_{i}$. Note that $h^{l;\psi}_{i}$ are with same shapes with $h^{l;\xi}_{i}$.
To alleviate the common class imbalance issue in SSMIS, a Difficulty-aware Re-weighting Strategy (DRS) is utilized to force the model to focus on the most difficult classes (i.e. the classes learned slower and with worse performances). The objective function of the supervised difficulty-aware training is defined as follow:
$\mathcal{L}_{diff}=\frac{1}{N_{L}}\frac{1}{K}\sum_{i=0}^{N_{L}}\sum_{k=0}^{K}w_{k}^{diff}\mathcal{L}_{DiceCE}(p^{l;\psi}_{i,k},y_{i,k})$ (3)
The overall difficulty-aware weight of $k^{th}$ class is defined as:
$w_{k}^{diff}=w_{\lambda_{k,e}}\cdot(d_{k,e})^{\alpha}$ (4)
$\centering d_{k,e}=\frac{{du_{k,e}}}{{dl_{k,e}}}\@add@centering$ (5)
$du_{k,e}=\sum^{e}_{e-\tau}\mathrm{min}(\triangle,0)\mathrm{ln}(\frac{\lambda_{k,e}}{\lambda_{k,e-1}}),$ (6)
$dl_{k,e}=\sum^{e}_{e-\tau}\mathrm{max}(\triangle,0)\mathrm{ln}(\frac{\lambda_{k,e}}{\lambda_{k,e-1}})$
where $\alpha$ is empirically set to $\frac{1}{5}$ in the experiments to alleviate outliers, $\lambda_{k}$ denotes the Dice score of $p^{l;\xi}$ of $k^{th}$ class in $e^{th}$ iteration and $\triangle=\lambda_{k,e}-\lambda_{k,e-1}$. $du_{k,e}$ and $dl_{k,e}$ denote classes not learned and learned after the tethte{th} iteration. $\tau$ is the number of accumulation iterations and is set to 50 empirically. More details can be seen in A&D [10].

### v3

# Diverse Teaching and Label Propagation for Generic Semi-Supervised Medical Image Segmentation

@@ Abstract
Limited annotation and domain shift severely limit medical image segmentation in clinical practice, necessitating unified approaches for semi-supervised segmentation (SSMIS), unsupervised domain adaptation (UMDA), and semi-supervised domain generalization (Semi-MDG). Existing methods are generally tailored to specific tasks and fail to generalize across these tasks due to error accumulation from noisy pseudo-labels and poor exploitation of data structure. In this paper, we employ a Diverse Teaching and Label Propagation Network (DTLP-Net) to boost the generic semi-supervised medical image segmentation. DTLP-Net involves a single student model and two diverse teacher models, where the first teacher decouples the labeled and unlabeled with a diffusion decoder, and the second is the mean-teacher model. Their outputs are then synergistically fused via an entropy-based ensemble to yield robust supervisory signals. The framework’s performance is further bolstered by three synergistic consistency strategy. First, a global-local consistency module leverages cross-set CutMix and masked image modeling to learn domain-invariant representations from both inter- and intra-sample contexts. Second, masked reconstruction on the feature level and knowledge distillation from the soft prediction is further utilized to alleviate the adverse impact of the noise present the hard pseudo labels. Finally, a voxel-level label propagation strategy explicitly models and enforces pairwise dependencies, enhancing spatial coherence in the final segmentation. We evaluate our proposed framework on five benchmark datasets for SSMIS, UMDA, and Semi-MDG tasks. The results showcase notable improvements compared to state-of-the-art methods across all five settings, indicating the potential of our framework to tackle more challenging SSL scenarios.
@@ I Introduction
Semi-supervised medical image segmentation (SSMIS) methods [1, 2, 3, 4, 5] has emerged as a cornerstone in medical image segmentation, offering a potent solution to the prohibitive cost and expertise required for the manual labeling process. However, the practical utility of conventional SSMIS is frequently undermined by a second, equally critical challenge: domain shift. This phenomenon, arising from variations in scanners, patient populations, and imaging protocols, causes severe performance degradation when models are deployed in real-world clinical settings. The confluence of these two challenges has given rise to more complex and realistic scenarios, including unsupervised domain adaptation (UMDA) [6, 7, 8], where the model must adapt from a labeled source domain to an entirely unlabeled target domain, and the even more demanding semi-supervised domain generalization (Semi-MDG)[9, 10]. The latter aims to train a model on a collection of partially labeled source domains to perform robustly on a completely unseen domain, representing a critical step toward true clinical applicability and robustness.
Despite their shared conceptual roots in learning from limited labels, these scenarios have historically been tackled in isolation. This fragmented approach has led to a proliferation of task-specific models that lack versatility; a state-of-the-art SSMIS method may fail entirely when faced with the domain shifts inherent to UMDA or Semi-MDG [11, 12]. While recent efforts like AIDE [11] and A&D [12] have pioneered a unified approach, their performance gains in cross-domain settings remain incremental. A&D [12] introduced an aggregating and decoupling approach to solve these three challenging tasks simultaneously; its advancements in tasks involving domain shift, e.g., UMDA and Semi-MDG, remain limited compared to prior art methods that are specifically designed for these tasks. We argue this is due to two fundamental, unresolved issues: (1) The generation of unreliable pseudo-labels, which become particularly noisy and damaging in the presence of domain shift, leading to error accumulation [13] and model collapse [14]. The main challenge in SSMIS lies in efficiently leveraging unlabeled images, while UMDA and Semi-MDG scenarios further require addressing the domain shift issues. All of these three tasks benefit from the reliable pseudo-label and model diversity [2]. (2) The under-utilization of intrinsic data structure, where voxel-level relationships and spatial context are not fully exploited to guide robust, domain-invariant feature learning. A&D [12] adopts the Diffusion-based V-Net to capture the domain-invariant features, which is not enough to learn the common knowledge of different domains.
To address this challenge, we propose a Diverse Teaching and Label Propagation Network (DTLP-Net) to boost the Generic Semi-Supervised Medical Image Segmentation, as shown in see Fig. 1. The core of our DTLP-Net is a sophisticated dual-teacher strategy engineered to generate reliable pseudo-labels, even in the presence of significant domain shift. Specifically, our approach builds on A&D [12], which features a shared diffusion encoder for domain-invariant feature extraction and a multi-decoder architecture (one Diffusion VNet, two V-Nets) to decouple the training process. For the unlabeled data stream, our diverse teaching mechanism generates high-quality supervision through two distinct teachers. The first teacher’s predictions are derived from the decoupled decoders, while the second is generated by a EMA mean-teacher [2]. We then employ an entropy-based ensembling strategy to intelligently fuse their outputs, encouraging the student to learn from both consistent and conflicting predictions, as illustrated in Fig. 1 (f). This cross-supervision compels the decoders to learn complementary features and correct each other’s mistakes, enhancing overall performance. Moreover, DTLP-Net incorporates three powerful consistency learning mechanisms. First, we leverage a cross-set Cut-Mix strategy and masked image modeling to enforce both global and local consistency, compelling the model to learn domain-agnostic structural features. Then, since the hard pseudo labels generated by the dual teachers still inevitably contain noise, especially in the scenario where there is class imbalance, we introduce additional supervision for the student. This includes masked reconstruction, which aligns the student’s predictions with the second mean-teacher’s logits, and knowledge distillation from the first teacher’s soft labels. Finally, we introduce a voxel-level label propagation module that explicitly models and enforces pairwise spatial similarities within an image, enhancing segmentation coherence and robustness. By synergizing diverse teaching with multi-level consistency, our framework effectively mitigates domain shifts and maximizes information extraction from unlabeled data. The overall contributions can be summarized as follows:
@@
  • We propose an effective and unified framework DTLP-Net to boost the Generic Semi-Supervised Medical Image Segmentation, including SSMIS, UMDA, and Semi-MDG tasks.
  • We introduce a novel dual-teacher, entropy-based ensemble strategy that generates high-quality pseudo-labels in the presence of domain shift. By explicitly leveraging teacher diversity and conflict, our method effectively reduces confirmation bias.
  • We design a hierarchical consistency mechanism that learns domain-invariant features across multiple scales. It enforces global-local structural priors through Cut-Mix and masking, while simultaneously ensuring fine-grained spatial coherence via a novel label propagation module. Moreover, we propose masked reconstruction and knowledge distillation to mitigate the adverse impact of the noise present in the hard pseudo labels.
@@
  • We evaluate our approach on five standard benchmark datasets, the experimental results show that our approach attains substantial enhancements across all three scenarios of SSMIS, UMDA, and Semi-MDG.
@@ II-A Semi-supervised medical image segmentation
In an effort to reduce the costs related to volume-level annotation while still ensuring high accuracy, diverse semi-supervised methods for medical image segmentation tasks have been developed. Drawing inspiration from semi-supervised image classification [15], techniques such as self-training and consistency regularization have emerged. In addition, the mean-teacher approach [16] and its extensions [17, 1, 18] have gained significant attention. For instance, BCP [1] advocates a straightforward Mean Teacher architecture to copy and paste annotated and unlabeled data bidirectionally, highlighting a key challenge in semi-supervised learning where the distributions of labeled and unlabeled data may diverge. Moreover, some works resort to the co-training paradigm, where multiple models are collaboratively trained by cooperatively training multiple models. CTCL [3] proposes a collaborative transformer-CNN learning for semi-supervised medical image segmentation. W2sPC [19] further incorporates weak-to-strong perturbation consistency and edge-aware contrastive representation, thus promoting the consistency of reliable regions among diverse predictions and facilitating the learning of class-discriminative representations. GapMatch [20] effectively links instance and model perturbations, thereby expanding the perturbation space. Moreover, it utilizes dual perturbation to enforce consistency regularization on the model. Different from these works that only consider one or two perturbations to enhance the model diversity, CMMT-Net [2] proposes a unified framework to generate reliable pseudo-labels and achieve consistency learning. In contrast to the aforementioned techniques, we adopt co-training by multi-decoder to enlarge the diversity of the models while designing multi-teachers to enhance the pseudo-labels quality. This simple yet effective approach improves the model’s generalization ability and segmentation accuracy.
@@ II-B Unsupervised Medical Domain Adaptation
Unsupervised Domain Adaptation (UDA) [6, 21] aims to solve the domain shifts by jointly training the model with labeled source domain data and unlabeled target domain data. In the field of Unsupervised medical Domain Adaptation (UMDA), one modality usually lacks any segmentation annotation, which has recently gained increasing attention since it offers an efficient way to compensate for limited medical image data. Existing UMDA approaches resort to various directions, e.g. semi-supervised learning [22], generative adversarial-based methods [23, 24], and contrastive learning [25], aiming to mitigate the adversarial influence caused by the severe domain shifts via image-level [6], feature-level [24] or both [21, 7]. Although with promising adaptation results, these methods fail to comprehensively exploit the unlabeled target domain information, which impedes generalization. Moreover, the output-level alignment across different domains still awaits exploration.
@@ II-C Semi-supervised Medical Domain Generalization
Semi-supervised domain generalization (SemiDG) is a more challenging setting, where the model does not use any information from the target domain, and the data in the source domains are only partially labeled. Compared with SSMIS and UMDA tasks, SemiDG necessitates the model to possess robust feature extraction capabilities and a high degree of generalizability. Existing SemiDG methods [26, 27] use various carefully designed strategies to solve the domain shifts, e.g., meta-learning [26], and Fourier transformation [9], which are not general and have unsatisfactory performance on the tasks such as UDA and SSL. Considering all of these three tasks involve learning from both labeled and unlabeled data, A&D [12] first try to unified with a generic framework handling all of them. However, despite the advancement they achieved, the improvement on SSMIS, UMDA and Semi-MDG tasks is marginal compared with task-tailored SOTAs.
@@ III Methodology
In the 3D Generic Semi-Supervised segmentation setting, the training data comprises $N_{L}$ labeled samples $D_{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{N_{L}}$ and $N_{U}$ unlabeled samples $D_{u}=\left\{x_{i}\right\}_{i=1}^{N_{U}}$, respectively, where $x_{i}\subset\mathbb{R}^{L\times W\times H}$ depicts the input volume with dimensions $L\times W\times H$, and $y_{i}\subset\left\{0,1\right\}^{Y\times L\times H\times W}$ represents the ground truth with $Y$ classes. The aim is to utilize the datasets $D_{l}$ and $D_{u}$ to train a model to generate meaningful semantic predictions. Our DTLP-Net consists of a shared encoder $\mathcal{E}(x^{l},x^{u};\xi)$ to learn domain-invariant features, and three decoders, e.g., V-Net Decoder $\mathcal{D}(x^{l};\psi)$, Diffusion Decoder $\mathcal{D}(x^{l},x^{u};\xi)$ and V-Net Decoder $\mathcal{D}(x^{u};\theta)$, where $\psi$, $\xi$, and $\theta$ are model parameters, as illustrated in Fig. 1. In addition, for $\mathcal{D}(x^{u};\theta)$ Decoder branch, we further obtain the mean-teacher Encoder $\mathcal{E}(x^{l},x^{u};\xi^{T})$ and Decoder $\mathcal{D}(x^{u};\theta^{T})$.
@@ III-A Preliminary baseline
$\small y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\epsilon\in\mathcal{N}(0,1)$ (1)
@@
$\small\mathcal{L}_{diff}=\frac{1}{N_{L}}\sum_{i=0}^{N_{L}}\mathcal{L}_{DiceCE}(p^{l;\xi}_{i},y_{i}),$ (2)
where $\mathcal{L}_{DiceCE}(x,y)=\frac{1}{2}[\mathcal{L}_{Dice}(x,y)+\mathcal{L}_{CE}(x,y)]$ is the combined dice and cross-entropy loss.
Supervised Training with $D(x^{l};\psi)$
For the difficulty-aware training flow, i.e., $D(x^{l};\psi)$ as decoder, the encoder only takes $x_{l}$ input to obtain the multi-scale features $h^{l;\psi}_{i}$. Note that $h^{l;\psi}_{i}$ are with the same shapes as $h^{l;\xi}_{i}$. The objective function of the supervised difficulty-aware training is defined as follows:
$\small\mathcal{L}_{sup}=\frac{1}{N_{L}}\frac{1}{K}\sum_{i=0}^{N_{L}}\sum_{k=0}^{K}\mathcal{L}_{DiceCE}(p^{l;\psi}_{i,k},y_{i,k})$ (3)